OpenStudy (anonymous):

Please help! Find the remaining parts of the right triangle having side a = 12 and side b = 16. Give the results properly rounded off. B = a. 53 degrees b. 37 degrees c. 49 degrees d. 41 degrees

OpenStudy (anonymous):

Here, you should be able to trivially obtain the hypotenuse (pythagorean theorem). Recall the law of sines also.

OpenStudy (anonymous):

 from the diagram, how can you find angle B?

OpenStudy (anonymous):

remember \[\tan \theta = opp/adj\]

OpenStudy (anonymous):

^ Easier solution

OpenStudy (anonymous):

\[\tan \theta = 16/12 = 1.33333\]